OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 6 and contains the point (1, 1)

OpenStudy (anonymous):

6(x-1)=y-1 y=6x-5

OpenStudy (alexwee123):

plug it into the point slope formula  where m=slope and x1 and y1 the point that you were given